### v1

# Efficient Egocentric Human Activity Recognition

Abstract
Egocentric video-based models capture rich semantic information and have demonstrated strong performance in human activity recognition (HAR). However, their high power consumption, privacy concerns, and dependence on lighting conditions limit their feasibility for continuous on-device recognition. In contrast, inertial measurement unit (IMU) sensors offer an energy-efficient and privacy-preserving alternative, yet they suffer from limited large-scale annotated datasets, leading to weaker generalization in downstream tasks. To bridge this gap, we propose COMODO, a cross-modal self-supervised distillation framework that transfers rich semantic knowledge from the video modality to the IMU modality without requiring labeled annotations. COMODO leverages a pretrained and frozen video encoder to construct a dynamic instance queue, aligning the feature distributions of video and IMU embeddings. By distilling knowledge from video representations, our approach enables the IMU encoder to inherit rich semantic information from video while preserving its efficiency for real-world applications. Experiments on multiple egocentric HAR datasets demonstrate that COMODO consistently improves downstream classification performance, achieving results comparable to or exceeding fully supervised fine-tuned models. Moreover, COMODO exhibits strong cross-dataset generalization. Benefiting from its simplicity, our method is also generally applicable to various video and time-series pre-trained models, offering the potential to leverage more powerful teacher and student foundation models in future research. The code is available at https://github.com/Breezelled/COMODO .
1 Introduction
Egocentric human activity recognition (HAR) has gained increasing attention due to its applications in healthcare, fitness tracking, and human-computer interaction. With the increasing availability of smart glasses, headsets, and next-generation wearable devices, these devices often integrate multiple sensors, including cameras and inertial measurement units (IMUs), enabling multimodal sensing for HAR.
Figure 1: Motivation: Egocentric videos provide rich semantic information but are impractical for continuous on-device recognition, while IMU sensors are lightweight and energy-efficient yet lack large-scale training data. To bridge this gap, we propose cross-modal, self-supervised distillation to enhance IMU representations by leveraging video knowledge.
Recent advances in large-scale video pretraining [40, 2, 27] have significantly enhanced video-based HAR. The availability of large-scale datasets [5, 3, 17, 18, 19] has fueled the success of deep learning models in this domain. Although the video modality achieves high accuracy in human activity recognition, it is often impractical for lightweight devices due to high power consumption, privacy concerns, varying lighting conditions, frequent interruptions, etc. Unlike cameras, IMUs operate independently of lighting conditions and visual occlusions while consuming significantly less power, making them ideal for always-on, on-device activity recognition in resource-constrained scenarios. This efficiency and robustness make IMUs a compelling alternative for HAR on lightweight devices. Despite these advantages, its available training data is significantly more limited. Unlike fields such as computer vision and natural language processing, which have seen significant advancements driven by large-scale datasets, human activity recognition using wearable sensors still relies on relatively small datasets, many of which are collected in controlled environments with scripted activities [47, 1, 35, 6]. A key challenge is that annotating IMU data is much more labor-intensive and costly than labeling images or text, as it often requires external reference signals and precise temporal alignment [21]. This lack of diverse real-world data makes it challenging for data-hungry deep learning models to learn robust representations, limiting their generalization ability in practical applications.
Given the limitations of IMU data, an ideal solution is to leverage the semantic richness of large-scale video training data and pretrained video models while maintaining the efficiency and privacy advantages of IMU sensors. However, there are several challenges to bridge the latent representation from the video domain to IMU domain. Firstly, we need an effective technique to effectively distill a superior video-domain knowledge to IMU domain. Secondly, we need a structured approach to align heterogeneous video and IMU modalities, given their distinct temporal resolutions and feature spaces. Thirdly, given the diversity of pretrained models and downstream tasks, the method needs to be general enough, adaptable to different video and IMU encoder pairs. Finally, the proposed method needs to be both data-efficient (self-supervised) and effectively leverage video representations to enhance IMU-based recognition.
We present CrOss-MOdal video-to-imu DistillatiOn (COMODO) for efficient egocentric HAR, a novel self-supervised learning framework designed to transfer knowledge from the video to IMU modality. Inspired by recent works in self-supervised learning [23], knowledge distillation [10, 29], and multimodal learning [31], our method maintains an instance queue to provide a consistent and large-scale distribution from teacher modality, facilitating the knowledge distillation between IMU and video features.
The main goal of cross-modal knowledge distillation is to leverage knowledge from data-rich modalities to enhance representation learning in data-scarce domains. We demonstrate that COMODO, by performing cross-modal self-supervised distillation, enables pretrained time-series model to inherit rich semantic information from video representations, significantly mitigating the challenges of data scarcity and annotation costs in IMU-based HAR. Our experiments across three benchmark datasets show that COMODO achieves performance comparable to or surpassing fully supervised fine-tuned time-series models, while maintaining the efficiency and privacy advantages of IMU sensors. These results suggest that cross-modal self-supervised learning can bridge the gap between high-resource and low-resource sensing modalities, providing a scalable and effective alternative for real-world wearable AI applications. Our contributions are summarized as follows: (1) Self-supervised Cross-modal Knowledge Transfer: We propose COMODO, a cross-modal self-supervised distillation framework that leverages pretrained video and time-series models enabling label-free knowledge transfer from a stronger modality (video) with richer training data to a weaker modality (IMU) with limited data. (2) A Self-supervised and Effective Cross-modal Queuing Mechanism: We introduce a cross-modal FIFO queue that maintains video embeddings as a stable and diverse reference distribution for IMU feature distillation, extending the instance queue distribution learning approach from single-modality to cross-modality. (3) Teacher-Student Model Agnostic: COMODO supports diverse video and time-series pretrained models, enabling flexible teacher-student configurations and future integration with stronger foundation models. (4) Cross-dataset Generalization: We demonstrate that COMODO maintains superior performance even when evaluated on unseen datasets, and more superior than fully supervised models, highlighting its robustness and generalizability for egocentric HAR tasks.
2 Related Works
Figure 2: Overview of our cross-modal self-supervised distillation framework. The video encoder is pretrained and kept frozen, while the IMU encoder, initialized from a pretrained time-series model, is trained by minimizing the cross-entropy loss between the similarity distributions of video and IMU embeddings, which are computed based on a continuously updated instance queue.
Human Activity Recognition (HAR) has been explored across both vision and sensor-based modalities, each offering unique advantages. Self-supervised learning (SSL) has emerged as a key approach in HAR for leveraging unlabeled sensor data, with contrastive learning effectively capturing generalizable features. [8, 9] introduced contrastive cross-modal learning frameworks for multi-sensor time-series, with [9] improving upon [8] by replacing explicit negative pairs with latent masking. [31] aligned IMU data with vision-language models [34] using contrastive loss, while [39] projected IMU embeddings into a video-based semantic space to enable zero-shot activity recognition. In contrast, our approach performs cross-modal self-supervised distribution distillation from video to IMU.
Knowledge distillation (KD) traditionally transfers knowledge from large teachers to smaller students. [38] proposes contrastive representation distillation to align teacher-student embeddings. Similarly, [37, 32] distill rich video representations into IMU models, with [37] needing video at inference and [32] requiring target data fine-tuning. [33] employs multimodal distillation, transferring knowledge from RGB, optical flow, and audio features to a single RGB model, demonstrating that multimodal supervision can improve egocentric HAR. [14] explore mutual distillation, where modality-specific models collaboratively enhance their representations in a supervised setting. For SSL, [10, 29] utilize self-supervised teachers to improve smaller models using the similarity distribution KD, facilitating the generalization of representation in visual and textual domains. [4] frames SSL as self-distillation with a momentum teacher. [43] shows that multimodal students can outperform unimodal teachers by learning richer cross-modal representations. Successful KD depends on the alignment of modality-general, task-relevant features [44].
Unlike [37, 32], which require either video input at inference or fine-tuning on the target dataset, our approach performs self-supervised cross-modal distillation, aligning sensor and video representations without labeled data. While prior self-supervised KD methods [10, 29, 4] focus on single-modal representation learning, our method explicitly transfers knowledge across heterogeneous modalities.
Knowledge transfer between modalities has also been explored using synthetic IMU generation. [25, 26] synthesized IMU signals from video and language models, demonstrating the potential for large-scale HAR by alleviating the need for extensive real-world IMU data collection. More recently, [28] aligned time-series sensor data with large language models, enabling HAR via natural language. These works focus on data synthesis and language-based IMU transfer. Multimodal HAR methods often integrate vision, IMU, and other sensors to improve recognition. [48] applied masked autoencoders for joint representation learning of IMU and video, while [13] introduced temporal attention to optimize multimodal feature integration. [30] proposed a self-supervised framework designed for multi-device HAR that learns robust sensor embeddings via spatial-temporal masking. [24] introduced hierarchical self-supervised pretraining to improve cross-dataset generalization, and [15] handled missing modality conditions through a combination of modality dropout training and cross-modal contrastive learning. To address domain adaptation in egocentric action recognition, [22] applied multimodal few-shot learning. These works reflect the trend of combining multimodal inputs (video, sensor, and language) to overcome data scarcity and domain shifts.
Compared to these works, we focus on cross-modal self-supervised distillation, leveraging video embeddings to enhance IMU representations without multimodal fusion at inference. Unlike prior contrastive learning methods that align sensor embeddings via instance-wise contrast, our approach distills structured distributional knowledge from video to IMU, enabling robust IMU-based HAR without requiring large-scale labeled datasets.
3 Methodology
As discussed in Section 1, while video-based models have demonstrated strong performance in HAR, they come with several practical limitations. On the other hand, IMU sensors are lightweight, power-efficient, and privacy-preserving, making them an attractive choice for real-world applications. However, IMU models lack access to large-scale labeled datasets, making it challenging to learn robust representations. To bridge this gap, we present our CrOss-MOdal self-supervised DistillatiOn (COMODO) that transfers knowledge from the video modality (which benefits from large-scale pretraining) to the IMU modality without requiring labeled data, as illustrated in Figure 2.
Contrastive SSL relies on a sufficiently large number of negative samples to learn discriminative representations effectively. However, maintaining an extremely large batch size is computationally prohibitive. To balance this trade-off, we adopt a FIFO queue approach [23]. This enables the model to maintain a large and diverse set of negative samples to learn from a much broader distribution of video embeddings while keeping the computational cost low.
We define our dataset as $D=\{(v_{i},x_{i})\}_{i=1}^{|D|}$, where each sample consists of a video window $v_{i}$ and its corresponding IMU window $x_{i}$ The dataset contains $|D|$ paired samples collected from egocentric wearable devices. Our objective is to learn an IMU encoder $f_{x}$ that maps an IMU window $x_{i}$ to a feature representation $z_{x}$ , such that the learned representations capture semantic information from the video modality without requiring labeled data.
3.1 Cross-modal Self-supervised Distillation
COMODO employs two separate pretrained encoders to process video and IMU data. Given a paired video-IMU window sample $(v_{i},x_{i})$ from the dataset, we extract the features using: (1) Video encoder $f_{v}$: A frozen pretrained video encoder that maps the video $v_{i}$ to a feature representation $f_{v}(v_{i})$. The video encoder remains frozen throughout training to leverage large-scale pretrained video representations. (2) IMU encoder $f_{x}$: A pretrained but trainable time-series encoder that maps the IMU signal $x_{i}$ to a feature representation $f_{x}(x_{i})$. Unlike the video encoder, the IMU encoder is continuously updated during training to distill knowledge from the video modality. Each encoder is followed by a modality-specific projection head: (1) Video projection head $f_{v}(v_{i})$ maps $h_{v}$ to a normalized embedding $z_{v}=h_{v}(f_{v}(v_{i}))$. (2) IMU projection head $f_{x}(x_{i})$ maps $h_{x}$ to a normalized embedding $z_{x}=h_{x}(f_{x}(x_{i}))$. To enforce cross-modal distribution soft alignment, both video and IMU embedding are L2-normalized.
The FIFO queue $Q$ maintains a large pool of video teacher embeddings. Given a new sample, the encoded video embedding $z_{vN}:=z_{v}$ of the current batch are enqueued $Q=\{z_{v1},z_{v2},\ldots,z_{vN}\}$, the IMU embedding $z_{x}$ and video embedding $z_{v}$ are then compared against all video embeddings stored in the queue to compute their similarity scores. To compute the similarity score distributions, we apply the softmax function over the inner product between the query embedding and all stored video embeddings:
$P_{v}(i|Q)=\frac{\exp(z_{v}\cdot z_{vi}/\tau^{v})}{\sum_{j=1}^{N}\exp(z_{v}\cdot z_{vj}/\tau^{v})},$ (1)
$P_{x}(i|Q)=\frac{\exp(z_{x}\cdot z_{vi}/\tau^{x})}{\sum_{j=1}^{N}\exp(z_{x}\cdot z_{vj}/\tau^{x})},$ (2)
[…]
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}P_{v}(i|Q)\log P_{x}(i|Q)$ (3)
After computing cross-entropy loss, the oldest batch of encoded videos $z_{v1}$ is dequeued to maintain a fixed queue size.
In COMODO, the video encoder $f_{v}$ is pretrained and kept frozen. Freezing $f_{v}$ ensures that the video embeddings stored in the FIFO queue $Q$ remain stable throughout training, providing a consistent reference for the IMU encoder to align with. Given that all embeddings are L2-normalized, the similarity score between $z_{v}$ and the most recent entry in the queue, $z_{vN}$, remains constant at 1 before softmax, as $z_{v}=z_{vN}$. This ensures that the highest similarity is always assigned to the latest video sample, while other entries in the queue provide a diverse set of negative samples for contrastive learning. By tuning the temperature factor $\tau^{v}$ and $\tau^{x}$, we can control the sharpness of the similarity distribution, adjusting the weight assigned to different video embeddings in the queue. The IMU encoder, trained to minimize cross-entropy with respect to the video similarity distribution, progressively aligns its representation with that of the video encoder, allowing it to leverage the rich semantic knowledge from the video modality.
Methods Self-supervised Ego4D EgoExo4D MMEA
Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5
DLinear [46] 10.60 25.27 36.74 32.20 59.91 78.16 24.01 37.73 47.18
Informer [49] 48.89 72.75 81.75 74.60 96.28 99.17 45.81 67.68 78.81
TimesNet [41] 44.33 68.02 77.79 69.20 93.44 98.04 50.30 74.24 83.99
MOMENT-small [16] 57.59 75.91 82.94 79.26 97.04 99.33 84.27 94.76 96.88
Mantis [11] 58.36 76.98 83.76 84.22 97.95 99.41 93.01 98.25 99.01
MOMENT-small [16] 39.70 64.47 75.32 68.14 93.55 98.43 83.66 95.52 97.42
Mantis [11] 47.49 71.63 81.24 76.47 96.98 99.21 90.96 98.56 99.39
IMU2CLIP [31] 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
COMODO (Ours) 59.13 78.79 85.65 84.92 98.28 99.59 92.48 99.01 99.77
Table 1: IMU-based Human Activity Recognition results. Our method, COMODO, achieves comparable or superior performance to fully supervised models, while consistently outperforming the previous state-of-the-art cross-modal self-supervised method, demonstrating its effectiveness in cross-modal transfer learning.
3.2 Inference
Once training is complete, only the IMU encoder $f_{x}$ is used for inference, while the video encoder $f_{v}$, projector networks $h_{v},h_{x}$, and the FIFO queue $Q$ are discarded. Given an IMU window $x$, the trained IMU encoder acts as a feature extractor, producing an embedding $z_{x}$. Following prior work on unsupervised representation learning [16, 45, 12], we evaluate the quality of the learned representations by training a Support Vector Machine (SVM) with an RBF kernel on the extracted features from the training subset, and then using it to predict the classes of the test subset. In this way, our representation learning distillation framework remains unsupervised, with class labels only used to train the SVM on the extracted features.
4 Experiments
4.1 Datasets and Preprocessing
To evaluate the effectiveness of COMODO, we used 3 publicly available egocentric HAR datasets: Ego4D [18], EgoExo4D [19], and MMEA [42]. The statistics of the datasets are summarized in Table 2.
Dataset Statistics Tra. Val. Tst.
Ego4D # Media files 1110 162 316
Total Media Durations 352.2h 51.9h 95.7h
[…]
# Activities 31 31 31
EgoExo4D # Media files 3495 501 998
Total Media Durations 153.8h 22.7h 42h
[…]
# Activities 8 8 8
MMEA # Media files 4553 653 1316
Total Media Durations 21.2h 3h 6.2h
[…]
# Activities 32 32 32
Table 2: Dataset statistics for Ego4D, EgoExo4D and MMEA.
MMEA provides raw sensor readings, including accelerometer and gyroscope data. Before use, we normalize the IMU data by dividing it by the provided sensitivity coefficients $R_{\text{acc}}=16384$, $R_{\text{gyro}}=16.4$ to obtain real acceleration values (in g) and gyroscope angular velocity (in ${}^{\circ}/s)$. The video data in MMEA is resized to 224 $\times$ 224 pixels, and its frame rate is maintained at 25 FPS. The IMU data are 25 Hz. Since this sampling rate is significantly lower than that of the other two datasets, it results in a relatively short sequence length, with an average of 418.57 and a median of 422. To address this, we apply left padding with an attention mask or resize the sequence length according to the requirements of different IMU student models.
For Ego4D and EgoExo4D, we follow the preprocessing steps of IMU2CLIP [31]. To ensure consistency across samples, we standardize the IMU data to a uniform sampling rate of 200 Hz. The videos are resized to 224 $\times$ 224 pixels and downsampled to 10 FPS. We divide each dataset into train, test, and validation sets following a 7:2:1 ratio. Since not all Ego4D videos include IMU recordings, we first exclude those without IMU data. After this initial selection, we further exclude activity categories that contain $<10$ video samples to ensure sufficient representation per class. Similar to IMU2CLIP, we segment both video and IMU data into non-overlapping, 5-second windows, discarding the final segment if its duration is $<$ 5 seconds. Given that the IMU data in these two datasets has a higher sampling rate, each 5-second segment results in a sequence length of $5\times 200=1000$ time steps. To accommodate different IMU student models, we either subsample or resize the sequences based on their specific input requirements.
4.2 Training
Setup. In the cross-modal self-supervised distillation stage, we train COMODO for 20 epochs with a learning rate of 3e-4 and a batch size of 128. In Equation 1 and 2, the video teacher tempature is set as $\tau^{v}$ = 0.1 and the IMU student temperature is $\tau^{x}$ = 0.05. All experiments are conducted on a single NVIDIA L40S GPU.
Video teacher. By default, we use the pre-trained TimeSformer-base (#params: 121.4M) [2] fine-tuned on Kinetics-400111https://huggingface.co/facebook/timesformer-base-finetuned-k400 [5] to leverage the rich knowledge from video modality. Following [7], we append an MLP projector $h_{v}$ at the end of the encoder after pooling. The hidden dimension of $h_{v}$ is 2048, the last feature dimension is 128. To ensure efficient training, all video teacher networks remain frozen throughout the distillation process. Instead of computing video features on the fly, we precompute and store all video embeddings in a pickle file, allowing direct retrieval during training to reduce computational overhead.
IMU student. We choose two lightweight time-series foundation models with significantly fewer learnable parameters as the IMU student network: (1) Mantis (#params: 8.11M) [11], the current state-of-the-art classification time-series foundation model which serves as the default choice, and (2) MOMENT-small (#params: 37.9M) [16]. To generate a unified IMU representation from multiple sensor channels, we apply concat pooling, where features from each IMU channel are processed separately and then concatenated. Specifically, given an IMU input sequence $x\in\mathbb{R}^{T\times C}$ with $C$ sensor channels and $T$ time steps, we apply a shared time-series model $f_{x}$ independently to each channel: $z_{c}=f_{x}(x_{c}),\forall c\in\{1,\dots,C\}$, where $x_{c}\in\mathbb{R}^{T}$ represents the input sequence for the $c$-th IMU channel, and $z_{c}\in\mathbb{R}^{D}$ is the corresponding feature representation where $D$ denotes the dimension. The final IMU embedding $z_{x}$ is obtained by concatenating all per-channel embeddings: $z_{x}=\{z_{1},z_{2},\dots,z_{C}\}\in\mathbb{R}^{C\times D}$. This approach ensures that the model captures independent channel-wise representations while maintaining structural consistency in the concatenated feature space. In Section 4.4.3, we present an ablation study showcasing results with different input pooling methods. Similar to the video teacher network, we add one additional MLP projector $h_{x}$.
4.3 Evaluation
To show the effectiveness of our proposed framework, we evaluate the representation quality of the IMU student encoder by comparing its performance against state-of-the-art time-series models and self-supervised learning methods on two downstream tasks. We use Accuracy @1, @3, and @5 to assess learned representations, as these metrics capture both the top-ranked prediction accuracy and the overall ranking quality within the predicted similarity distribution.
4.3.1 IMU-based Human Activity Recognition
This task aims to classify human activities from IMU sensor data for efficient and reliable recognition on resource-constrained wearables. Effective IMU-based HAR models should capture motion semantics and generalize robustly across diverse real-world sensor readings. We benchmark the classification on three egocentric human activity recognition datasets. Table 1 categorizes methods into supervised and Self-supervised settings:
Supervised methods. We include state-of-the-art time-series models as our supervised baselines: DLinear [46], Informer [49], TimesNet [41], MOMENT [16], and Mantis [11]. These models are fine-tuned on labeled IMU data.
Self-supervised methods. We evaluate pretrained time-series foundation models as our self-supervised baselines: MOMENT-small [16] and Mantis [11], which are assessed in a zero-shot manner without fine-tuning or distillation on the target IMU dataset. Additionally, we include IMU2CLIP [31] as our contrastive learning baseline, which leverages cross-modal self-supervised knowledge transfer to allow the IMU student model to acquire knowledge from the video modality, similar to COMODO. For these self-supervised methods, we follow the evaluation protocol in Section 3.2. To ensure a fair comparison, we re-implemented IMU2CLIP, preserving its core contrastive learning framework, and replaced the original CLIP-based teacher with our video teacher network and substituted IMU2CLIP’s student encoder with our IMU student network.
Results. As shown in Table 1, COMODO achieves the highest accuracy on most metrics across all three datasets. On Ego4D, it outperforms the state-of-the-art cross-modal self-supervised baseline IMU2CLIP by 4.3% in Acc@1 and matches the best supervised fine-tuned time-series baseline Mantis with a 0.77% improvement. On EgoExo4D, it outperforms IMU2CLIP by 5.46% in Acc@1 and exceeds supervised fine-tuned Mantis by 0.7%. On MMEA, COMODO surpasses IMU2CLIP by 1.07% in Acc@1 while closely matching the best supervised method, falling short by only 0.53%. Notably, COMODO achieves the highest scores in Acc@3 and Acc@5, indicating that it captures richer semantic representations, leading to a better overall prediction distribution. These results highlight the effectiveness of COMODO in leveraging cross-modal knowledge transfer to enhance IMU-based HAR. Furthermore, its ability to match or even surpass fully supervised fine-tuned models demonstrates its potential as a scalable and label-efficient solution for HAR and other related tasks.
Train Test Method Acc@1 Acc@3 Acc@5
EgoExo4D Ego4D Mantis 52.58 75.55 83.83
IMU2CLIP 51.75 74.44 83.05
COMODO 56.62 77.25 84.79
Ego4D EgoExo4D Mantis 80.79 97.82 99.53
IMU2CLIP 79.70 97.68 99.45
COMODO 82.54 98.02 99.56
Ego4D MMEA Mantis 91.72 98.48 99.32
IMU2CLIP 92.78 98.78 99.47
COMODO 94.22 98.78 99.62
EgoExo4D MMEA Mantis 92.25 98.10 99.16
IMU2CLIP 92.78 98.93 99.70
COMODO 94.83 99.47 99.70
Table 3: Cross-Dataset Generalization results showing transferability to unseen data. $\dagger$ indicates the model was supervised fine-tuned on the train dataset.
IMU Model Video Model Methods Ego4D EgoExo4D MMEA
Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5 Acc@1 Acc@3 Acc@5
MOMENTS VideoMAEB-k400 IMU2CLIP 41.39 65.82 76.45 69.77 94.61 98.81 78.42 91.19 94.60
COMODO 57.74 76.66 83.79 81.60 97.68 99.48 79.48 91.19 94.45
MOMENTS TimeSformerB-k400 IMU2CLIP 43.59 68.17 78.06 70.73 95.15 98.98 78.80 92.17 95.14
COMODO 57.51 76.49 83.75 81.84 97.93 99.54 81.16 92.25 95.36
Mantis VideoMAEB-k400 IMU2CLIP 54.43 75.92 83.89 79.09 97.57 99.44 90.88 98.63 99.47
COMODO 59.00 78.98 85.95 84.46 98.31 99.56 93.54 99.24 99.77
Mantis TimeSformerB-k400 IMU2CLIP 54.83 76.04 84.03 79.46 97.68 99.44 91.41 98.71 99.47
COMODO 59.13 78.79 85.65 84.92 98.28 99.59 92.48 99.01 99.77
Table 4: Result for different video teacher IMU student model combinations on Ego4D, EgoExo4D, and MMEA.
4.3.2 Cross-Dataset Generalization
The goal of this task is to measure the capability of a model to generalize its learned representations to other datasets that differ significantly from its training set. Strong cross-dataset generalization is critical for deploying HAR models in diverse and realistic environments, where variations in sensor setups, activities, and user behaviors are common. To evaluate this capability, we train models on one dataset and test them on another without additional fine-tuning, following the evaluation protocol in Section 3.2 to assess zero-shot generalization. We experiment with the following setups: (1) EgoExo4D $\rightarrow$ Ego4D, (2) Ego4D $\rightarrow$ EgoExo4D, (3) Ego4D $\rightarrow$ MMEA, and (4) EgoExo4D $\rightarrow$ MMEA.
Results. Table 3 presents the results of our cross-dataset generalization experiments. Notably, COMODO improves over IMU2CLIP and supervised fine-tuned Mantis baselines in all cases. In the EgoExo4D → Ego4D setup, COMODO outperforms IMU2CLIP by 4.87% in Acc@1 while also surpassing the supervised fine-tuned Mantis by 4.04%. When transferring from Ego4D to EgoExo4D, COMODO achieves the highest Acc@1 of 82.54%, improving upon the Mantis baseline by 1.75%. On the MMEA dataset, COMODO maintains its superiority, surpassing Mantis by 2.5% and 2.58% when trained on Ego4D and EgoExo4D, respectively. These results underscore COMODO’s robustness in adapting to datasets with significantly different distributions, further demonstrating the effectiveness of cross-modal self-supervised knowledge transfer in enhancing IMU representation learning.
4.4 Ablation Study
4.4.1 Different Teacher & Student Networks
To evaluate the generalizability of COMODO across different models, we conduct experiments using two IMU student models (Mantis [11] and MOMENT-small [16]) and two video teacher models (VideoMAE-Base-k400222https://huggingface.co/MCG-NJU/videomae-base-finetuned-kinetics [40] and TimeSformer-Base-k400 [2]). The results in Table 4 demonstrate that COMODO consistently improves performance across the vast majority of teacher-student configurations, validating its robustness in different settings.
We observe that TimeSformer-Base-k400 generally outperforms VideoMAE-Base-k400, suggesting that a stronger spatio-temporal feature extractor benefits IMU representation learning. Similarly, Mantis surpasses MOMENT-small as an IMU student, aligning with its superior standalone performance in Table 1. Notably, COMODO achieves significant gains across all networks, demonstrating its teacher-student model agnosticism and adaptability to various video and time-series models.
Figure 3: Impact of queue size on accuracy across datasets.
4.4.2 FIFO Queue Size
The optimal queue size $|Q|$ is dataset-specific, as we observe that different datasets benefit from different queue configurations. Our ablation study (Figure 3) reveals that increasing queue size does not always yield better results. Instead, the relationship between queue size and performance is influenced by the characteristics of the dataset:
Ego4D ($|Q|$ = 16384): As a large-scale dataset with diverse activities (31 classes), increasing the queue size generally improves performance by providing a richer set of negative samples for contrastive learning. However, we observed that the largest queue size tested (65536) does not yield the best performance. Prior research [29] also suggests that the optimal queue size is not necessarily the largest available, and an overly large queue may not always lead to improved results.
EgoExo4D ($|Q|$ = 1024): Empirical results indicate that increasing the queue size initially improves performance, but beyond $|Q|$ = 1024, further increasing the queue size leads to a decline in accuracy (Figure 3). Unlike Ego4D and MMEA, EgoExo4D has only 8 activity classes, leading to highly concentrated feature distributions. When the queue size is large, excessive redundant negatives from the same limited set of categories create sharp similarity spikes in the distribution. This makes it harder for the student model to align with the teacher’s representation. Prior work [20] suggests that KL divergence is easier to minimize when the distribution has a single dominant mode, whereas in the distributions of multiple distinct modes, forward KL forces the model to cover all modes, potentially assigning probability mass to less relevant regions. A moderate queue size helps regulate the number of distinct similarity peaks, stabilizing training and improving teacher-student alignment.
MMEA ($|Q|$ = 2048): This dataset contains 32 activity classes but has significantly fewer IMU samples compared to the other datasets. Given the limited number of samples, a very large queue is impractical. Instead, selecting a queue size that is proportionally appropriate for the dataset ensures that each mini-batch includes sufficiently diverse negatives. Our experiments show that a queue size of 2048 achieves the best performance on MMEA (Figure 3). Based on this observation, we choose 2048 as the queue size, which is relatively large for this dataset but helps maintain a balance between negative sample diversity and dataset constraints.
Train Method Acc@1 Acc@3 Acc@5
Ego4D Mean 55.69 75.21 82.92
Concat 57.51 76.49 83.75
EgoExo4D Mean 80.86 97.48 99.42
Concat 81.84 97.93 99.54
MMEA Mean 74.32 88.91 93.09
Concat 81.16 92.25 95.36
Table 5: Comparison of mean pooling and concat pooling for IMU representation learning across different datasets.
4.4.3 Pooling Methods
We compare mean pooling and concat pooling to evaluate their impact on IMU representation learning. Mean pooling compresses features by averaging across all sensor channels, while concat pooling preserves channel-wise distinctions by concatenating the features from each channel. For this ablation study, we use MOMENT-small as the IMU student model, as it provides built-in support for both pooling methods. In contrast, Mantis is designed with concat pooling as its default architecture, making it less straightforward to modify for this comparison.
Table 5 shows that concat pooling consistently outperforms mean pooling across all datasets. On Ego4D, it improves Acc@1 by 1.82%, while on EgoExo4D, the gain is 0.98%. The most significant improvement is observed on MMEA, where Acc@1 increases by 6.84%, indicating that maintaining per-channel information is especially beneficial for this dataset. These results suggest that mean pooling overly compresses features, leading to information loss, while concat pooling retains richer representations, enhancing classification accuracy. Given its superior performance across all datasets, we adopt concat pooling as the default method in our framework, as discussed in Section 4.2.
Figure 4: Accuracy of distillation methods across datasets.
4.4.4 Distillation Methods
To assess the effectiveness of our distillation strategy, we compare our strategy against two other distillation methods.
(1) L2 Distance Minimization [36]: L2 distance minimization directly aligns the student and teacher representations by minimizing the squared L2 distance between the IMU student embedding $z_{xi}$ and the video teacher embedding $z_{vi}$ for an identical paired video-IMU sample ($v_{i},x_{i}$). This method assumes that the feature spaces of different modalities are inherently aligned. As a result, L2 minimization can lead to suboptimal feature alignment.
(2) InfoNCE-based Contrastive Learning: InfoNCE optimizes the student encoder by maximizing the similarity between positive video-IMU pairs while minimizing similarity with randomly sampled negative instances. While effective in learning discriminative features, this method treats all negative samples equally, potentially misclassifying semantically similar negatives as incorrect, which limits its adaptability in cross-modal distillation.
(3) COMODO (ours): COMODO extends contrastive learning by distilling the similarity distribution rather than relying solely on instance-wise alignment. Unlike InfoNCE, which assumes all negatives are equally dissimilar, COMODO preserves the structure of the teacher’s similarity space, leading to more robust and generalizable feature representations. Figure 4 shows that COMODO consistently outperforms L2 and InfoNCE across all datasets, demonstrating the effectiveness of similarity distribution distillation in cross-modal representation learning.
5 Conclusion
In this work, we introduced COMODO, a cross-modal self-supervised distillation framework that leverages pretrained video representations to enhance IMU-based egocentric HAR. By distilling rich semantic knowledge from a frozen video encoder to an IMU encoder through a dynamic instance queue, COMODO effectively mitigates the challenges posed by limited sensor data and costly annotations. Extensive experiments across multiple egocentric HAR benchmarks demonstrate that COMODO achieves results comparable to or even surpassing fully supervised baselines. Notably, our method exhibits strong cross-dataset generalization, underscoring its robustness and practical utility. COMODO’s adaptability to various pretrained video and sensor models offers a viable framework for further investigating cross-modal knowledge transfer from visual models to resource-constrained sensors, enhancing future research in visually-guided, sensor-based HAR.
6 Acknowledgment
This research includes computations using the Wolfpack computational cluster, supported by the School of Computer Science and Engineering at UNSW Sydney. We also acknowledge support from the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005).